# TIPS: Text-Image Pretraining with Spatial awareness

### v1

@@ Abstract
While image-text representation learning has become very popular in recent years, existing models tend to lack spatial awareness and have limited direct applicability for dense understanding tasks. For this reason, self-supervised image-only pretraining is still the go-to method for many dense vision applications (e.g. depth estimation, semantic segmentation), despite the lack of explicit supervisory signals. In this paper, we close this gap between image-text and self-supervised learning, by proposing a novel general-purpose image-text model, which can be effectively used off the shelf for dense and global vision tasks. Our method, which we refer to as Text-Image Pretraining with Spatial awareness (TIPS), leverages two simple and effective insights. First, on textual supervision: we reveal that replacing noisy web image captions by synthetically generated textual descriptions boosts dense understanding performance significantly, due to a much richer signal for learning spatially aware representations. We propose an adapted training method that combines noisy and synthetic captions, resulting in improvements across both dense and global understanding tasks. Second, on the learning technique: we propose to combine contrastive image-text learning with self-supervised masked image modeling, to encourage spatial coherence, unlocking substantial enhancements for downstream applications. Building on these two ideas, we scale our model using the transformer architecture, trained on a curated set of public images. Our experiments are conducted on $8$ tasks involving $16$ datasets in total, demonstrating strong off-the-shelf performance on both dense and global understanding, for several image-only and image-text tasks.
@@ 2 Related Work
Image-text learning for dense understanding tasks. While the above-mentioned existing image-text learning approaches lead to powerful representations, they have not demonstrated clear benefits for dense image prediction tasks. As a consequence, today the models learned with self-supervised techniques are preferred for these cases: for example, the recent DepthAnything (Yang et al., 2024) model is built on top of the self-supervised DINOv2 features, even though weakly-supervised image backbones are widely available. Similarly, Tong et al. (2024) have demonstrated shortcomings of CLIP-style models and incorporated DINOv2 to enhance the visual grounding of multimodal models. However, recent work has adapted image-text learning for dense prediction, e.g. for open-vocabulary detection (Kim et al., 2023; Minderer et al., 2022; Rao et al., 2022) and segmentation (Mukhoti et al., 2023; Wu et al., 2024; Wysoczanska et al., 2024). SLIP (Mu et al., 2021) combines an adapted SimCLR (Chen et al., 2020) self-supervised objective with CLIP for classification tasks. Closer to our work, the very recent SILC method (Naeem et al., 2024) proposes to combine contrastive image-text training with self-distillation. Different from our goals, most of these methods are specifically tailored towards improving vision-language tasks that require spatial understanding and do not aim to learn a general-purpose vision encoder. Besides, most of these do not incorporate dense objectives during pretraining, and may require additional fine-tuning stages or dense supervision, which can be costly. SILC proposes to incorporate self-supervised losses into image-text training, but only employs self-distillation; in contrast, our TIPS technique goes beyond to employ a more sophisticated self-supervised approach leveraging masked image modeling with an online tokenizer, which we show to boost performance significantly for dense image prediction. FLIP (Li et al., 2023) proposed to combine contrastive learning with masking, but without any reconstruction loss, aiming only at efficient language-image training. Altogether, our approach demonstrates for the first time a way to learn image-text models whose vision representations rival those of self-supervised approaches in dense image understanding tasks.
@@ 3.1 Enhancing Weak Supervision with Synthetic Image Captions
A limitation of standard image-text learning using large-scale web data is the quality of the captions, which are noisy and may not accurately describe images. An example is shown in Fig. 3 (top), where the words “for sale dealership \$30k” are not describing the image contents. While this may hinder model learning, the caption still captures the main object (‘2007 Cadillac Escalade’).
@@
We employ off-the-shelf, publicly available models which can caption images effectively: given the image $I$, we generate text $\hat{T}$. In particular, we leverage captioning models which tend to generate accurate and high-level image descriptions – an example is given in Fig. 3 (bottom). Note the use of the preposition “in front of", which indicates the spatial arrangement of the scene, the description of the background (“building") and the color of the object (“black"), all combined providing rich signals for dense image representation learning.
@@
Dual image-text embedding. We strive to leverage relevant information from both captions, and thus propose to modify the vision transformer $f$ to learn from them, in an approach we call “dual embedding”. We insert an additional [CLS] token in the model, to be used for learning with the synthetic caption, obtaining an additional global embedding $\mathbf{\hat{e}}^{g}$. At training time, we feed both $T$ and $\hat{T}$ into the text encoder, to obtain their text embeddings $\mathbf{e}^{t}$ and $\mathbf{\hat{e}}^{t}$. In addition to the $\mathcal{L}_{CLIP}$ loss between $\mathbf{e}^{g}$ and $\mathbf{e}^{t}$, we compute ^$\hat{\mathcal{L}}_{CLIP}$ between $\mathbf{\hat{e}}^{g}$ and $\mathbf{\hat{e}}^{t}$. This introduces flexibility in the model to learn an object-centric image embedding in $\mathbf{e}^{g}$, and a more spatially-aware image embedding in $\mathbf{\hat{e}}^{g}$. Both back-propagate into the dense feature maps to learn improved patch embeddings $\{\mathbf{e}_{n}\}_{n=1}^{N}$. At inference time, the model can have access to both types of global image embeddings, and the one to use may be decided based on the downstream task: generally spatially aware tasks will use $\mathbf{\hat{e}}^{g}$ while object-centric ones will employ $\mathbf{e}^{g}$.
@@ 3.2 Integrating Self-Distillation and Masking to Boost Image Features
Discussion. Our method builds on representation learning ideas from the weakly and self-supervised literature, and to the best of our knowledge is the first to demonstrate that combining contrastive image-text learning with self-distillation and masked image modeling leads to improvements across many tasks, indicating positive synergies between these objectives. The closest existing technique is SILC (Naeem et al., 2024), which combined CLIP with self-distillation. As we show in experimental ablations, though, the addition of the masked image loss substantially improves performance across dense tasks, being critical for downstream applications. We note some key differences compared to previous methods. Given that we use a CLIP loss, the self-supervised components can be simplified, compared to the original formulations in DINO (Caron et al., 2021) and iBOT (Zhou et al., 2022). A major difference is that we use a single global “crop", instead of two in DINO, iBOT and SILC, substantially increasing throughput by $25\%$. In contrast to many self-supervised methods, we use comparatively simple data augmentations: the local crop is just a random crop of the original image, and the global crop is a larger random crop with horizontal flips. This is similar to Assran et al. (2023); Moutakanni et al. (2024) who argue that complex augmentations may not be necessary for representation learning. Finally, our masking approach is simply random, in contrast to blockwise in iBOT.
@@ 3.3 Scaling TIPS
Data. We leverage the WebLI dataset (Chen et al., 2023), which is a large and noisy web dataset of public images and associated alt-text containing $10$B image-text pairs. We filter the dataset in successive rounds in order to enhance its quality for model training, similar to previous work in language (Gunasekar et al., 2023; Wenzek et al., 2020) and vision (Oquab et al., 2024; Parthasarathy et al., 2023). This is critical for our model, since it is intended for off-the-shelf use in many downstream applications. First, similar to Schuhmann et al. (2022), we filter the image-text pairs based on their contents, by discarding those whose image-text similarities are low, as computed by a pretrained alignment model. Second, we filtered the resulting dataset to only keep pairs with English captions. These two initial steps result in a dataset of $1.7$B images. Finally, we follow a similar curation process as previous work (Oquab et al., 2024; Parthasarathy et al., 2023) and select images that are similar enough to those in curated datasets, leveraging a pretrained model to compute image embeddings; further details on the curated datasets and filtering strategies are detailed in Appendix A.3. Note that we also remove near-duplicate images from our dataset if they appeared in any of the evaluation datasets used in this paper. This process generates our main curated pretraining dataset, containing $116$M image-text pairs in total.
@@ 4.1 Experimental Setup
Evaluation datasets and protocols. Our models are evaluated on a suite of $8$ tasks involving $16$ datasets in total, comprising images-only or images-and-text tasks. We assess the quality of the learned representations thoroughly in a wide range of conditions, covering indoor/outdoor scenes and object-centric captures. Note that, in all evaluations, our image-text representations are kept frozen, since our goal is to assess their applicability as off-the-shelf feature extractors. We evaluate $3$ dense prediction tasks, $2$ holistic global image understanding tasks and $3$ multimodal retrieval tasks. We introduce the tasks below, and provide further details about their evaluation protocols in the appendix (Section A.4).
Semantic segmentation is a dense task evaluated on the PASCAL (Everingham et al., 2010) and ADE20k (Zhou et al., 2017) datasets, using mean Intersection over Union (mIoU). We use a simple linear probe setup similar to (Oquab et al., 2024), where classes are predicted from the spatial features.
@@
Zero-shot classification is conducted on ImageNet-$1$K by retrieving the class text embedding closest to the each test image’s embedding, following  (Radford et al., 2021), and using top-1 accuracy.
Additional training dataset details. As discussed in Sec. 3.3, we use images from a set of curated datasets as queries for mining among a large pool of web images. Following the steps adopted by DINOv2 (Oquab et al., 2024), we use the training sets of some of our evaluation datasets as the curated queries, details provided in the appendix. This leads to a web-based training dataset with $116$M image-text pairs. Additionally, for the scaling ViT-g experiment, we add the training set of the Mapillary SLS dataset (Warburg et al., 2020) as-is to our training set to compensate for the lack of street-level imagery in web images, and in the absence of any alt-text we use the generated synthetic caption for training both CLS tokens. This increases the number of images in our training set to $117$M in total. A similar procedure is conducted by DINOv2 for their LVD-$142$M dataset.
Implementation details. For all our models, we use $1$ global crop at resolution $224$ and $M=6$ local crops at resolution $98$. We train the ViT-B models for $70$ epochs at batch size $16$k, which takes $4$ days on $256$ TPUv3 chips. For the larger ViT-g model we train for $15$ epochs at batch size $16$k, which takes $2$ days on $512$ TPUv5 chips, and results in our low-res model (TIPS-g/14 LR). For our high-res variant (TIPS-g/14 HR), we add an additional finetuning stage with global crops at resolution $448$ and local crops at resolution $140$, which runs for $0.1$ epochs at batch size $4$k. We use only random resize crops and horizontal flips as image augmentations. Other details can be found in Sec. A.2.
Compared techniques. We strive to provide a large number of comparisons against recent work. For each existing model family, we compare against the largest instantiation up to ViT sizes “g” or “G”, at about $1.8$B parameters or less in the image encoder. We benchmark our method against a wide range of methods, from the self-supervised, weakly-supervised and supervised literature. All methods are used off-the-shelf, with frozen weights, for fair comparisons. As self-supervised methods, we compare against DINO (Caron et al., 2021), MAE (He et al., 2022), iBOT (Zhou et al., 2022) and DINOv2 (Oquab et al., 2024). As weakly-supervised methods, we compare against CLIP (Radford et al., 2021), OpenCLIP (Cherti et al., 2023), SigLIP (Zhai et al., 2023), SILC (Naeem et al., 2024) and EVA-CLIP (Sun et al., 2023). As a supervised method, we benchmark against the ViT-g trained on JFT-$3$B, as per (Zhai et al., 2022).
4.2 Results
@@
(PASCAL) (NYUv2) (ImageNet) (Flickr) (Flickr)
@@
Table 2: Image-only evaluations for dense and global prediction tasks. Experiments using the largest backbone available for each model variant, comparing recent self-supervised and image-text models against TIPS. Rows that are highlighted refer to image-only models that are very good at dense prediction tasks, but are by nature unable to perform text-related tasks. We also highlight the best and second-best number of each column. TIPS achieves the best or second-best performance in $7$ out of $9$ evaluations.
@@
Part (B) of the table ablates the contribution of enhanced textual supervision. Simply replacing the web captions by PaliGemma-generated ones improves segmentation by $10.1$ percentage points and reduces depth RMSE by 0.076, which are big positive gains. This shows the potential of synthetic captions for dense understanding with image-text models. However, at the same time global tasks show significant regressions, with KNN classification loss of $6.9$ points. But the CLIP performance can be improved in all tasks by combining the web and synthetic captions: using our dual embedding approach, we achieve large gains across the board. We also compare our dual approach against two other caption combination options, inspired by the ones proposed by Fan et al. (2023): “sampled”, where either the web or the synthetic caption is chosen at random; or “multi-text”, where both captions are matched against the same image embedding. Our dual approach performs better than other caption combinations in $3$ out $5$ cases and achieves competitive results on the other $2$, which indicates its effectiveness.
@@
Application: Single-image to 3D. Modern large reconstruction models rely on high-quality pre-trained image encoders to produce image tokens for an encoder/decoder transformer (Hong et al., 2024; Wang et al., 2024). For example, LRM (Hong et al., 2024) predicts parameters of a neural rendering model from the image features of a single input image. The authors choose the ViT-based DINO encoder over more semantic-aware ones (such as CLIP) due to its knowledge of structural and texture information necessary for 3D tasks.
@@
Table 3: Image-text evaluations for multimodal retrieval and zero-shot classification, where TIPS outperforms others in $6$ out of $7$ cases. Note that we compare solely against weakly-supervised methods, since self-supervised ones are not naturally aligned with language. We highlight the best and second-best number of each column.
@@
In order to better understand our model’s capabilities for neural 3D reconstruction, we evaluate TIPS performance in the LRM framework and compare DINO-B/16 to an equivalently-sized TIPS-B/14. We opt to use DINO-B/16 to follow after the original paper’s implementation. Single-image to $3$D results on the Objaverse (Deitke et al., 2023) dataset are presented in Tab. 4, showing that TIPS outperforms DINO as an image encoder for large reconstruction models, with enhanced novel view synthesis capabilities ($0.62$ increase in PSNR). Additional qualitative results can be found in A.4.4.
@@
Figure 4: Qualitative results for dense prediction tasks. For a given image (first column), we illustrate the principal components of the predicted spatial features (column 2) . Depth (column 3) and normals (column 4) are trained on NYUD, and for semantic segmentation (last column) we used the model trained on ADE20k. All dense tasks used the DPT decoder, while keeping the image encoder weights frozen. More qualitative results can be found in Section A.5.

### v2

@@ Abstract
While image-text representation learning has become very popular in recent years, existing models tend to lack spatial awareness and have limited direct applicability for dense understanding tasks. For this reason, self-supervised image-only pretraining is still the go-to method for many dense vision applications (e.g. depth estimation, semantic segmentation), despite the lack of explicit supervisory signals. In this paper, we close this gap between image-text and self-supervised learning, by proposing a novel general-purpose image-text model, which can be effectively used off the shelf for dense and global vision tasks. Our method, which we refer to as Text-Image Pretraining with Spatial awareness (TIPS), leverages two simple and effective insights. First, on textual supervision: we reveal that replacing noisy web image captions by synthetically generated textual descriptions boosts dense understanding performance significantly, due to a much richer signal for learning spatially aware representations. We propose an adapted training method that combines noisy and synthetic captions, resulting in improvements across both dense and global understanding tasks. Second, on the learning technique: we propose to combine contrastive image-text learning with self-supervised masked image modeling, to encourage spatial coherence, unlocking substantial enhancements for downstream applications. Building on these two ideas, we scale our model using the transformer architecture, trained on a curated set of public images. Our experiments are conducted on $8$ tasks involving $16$ datasets in total, demonstrating strong off-the-shelf performance on both dense and global understanding, for several image-only and image-text tasks. Code and models are released at https://github.com/google-deepmind/tips.
@@ 2 Related Work
Image-text learning for dense understanding tasks. While the above-mentioned existing image-text learning approaches lead to powerful representations, they have not demonstrated clear benefits for dense image prediction tasks. As a consequence, today the models learned with self-supervised techniques are preferred for these cases: for example, the recent DepthAnything (Yang et al., 2024) model is built on top of the self-supervised DINOv2 features, even though weakly-supervised image backbones are widely available. Similarly, Tong et al. (2024) have demonstrated shortcomings of CLIP-style models and incorporated DINOv2 to enhance the visual grounding of multimodal models. However, recent work has adapted image-text learning for dense prediction, e.g. for open-vocabulary detection (Kim et al., 2023; Minderer et al., 2022; Rao et al., 2022) and segmentation (Mukhoti et al., 2023; Wu et al., 2024; Wysoczanska et al., 2024). SLIP (Mu et al., 2021) combines an adapted SimCLR (Chen et al., 2020) self-supervised objective with CLIP for classification tasks. Closer to our work, MaskCLIP (Dong et al., 2023) leverages masked image modeling with contrastive learning, and the very recent SILC method (Naeem et al., 2024) combines contrastive image-text training with self-distillation. Different from our goals, most of these methods are specifically tailored towards improving vision-language tasks that require spatial understanding and do not aim to learn a general-purpose vision encoder. Besides, most of these do not incorporate dense objectives during pretraining, and may require additional fine-tuning stages or dense supervision, which can be costly. MaskCLIP and SILC propose to incorporate self-supervised losses into image-text training, but only employ either masked image modeling or self-distillation, respectively; in contrast, our TIPS technique goes beyond to combine both, which we show to boost performance significantly for dense image prediction. FLIP (Li et al., 2023) proposed to combine contrastive learning with masking, but without any reconstruction loss, aiming only at efficient language-image training. Altogether, our approach demonstrates for the first time a way to learn image-text models whose vision representations rival those of self-supervised approaches in dense image understanding tasks.
@@ 3.1 Enhancing Weak Supervision with Synthetic Image Captions
A limitation of standard image-text learning using large-scale web data is the quality of the captions, which are noisy and may not accurately describe images. An example is shown in Fig. 3 (top), where the words “for sale dealership \$30k" are not describing the image contents. While this may hinder model learning, the caption still captures the main object (‘2007 Cadillac Escalade’).
@@
We employ off-the-shelf, publicly available models which can caption images effectively: given the image $I$, we generate text $\hat{T}$. In particular, we leverage captioning models which tend to generate accurate and high-level image descriptions – an example is given in Fig. 3 (bottom). Note the use of the preposition "in front of", which indicates the spatial arrangement of the scene, the description of the background ("building") and the color of the object ("black"), all combined providing rich signals for dense image representation learning.
@@
Dual image-text embedding. We strive to leverage relevant information from both captions, and thus propose to modify the vision transformer $f$ to learn from them, in an approach we call "dual embedding". We insert an additional [CLS] token in the model, to be used for learning with the synthetic caption, obtaining an additional global embedding $\mathbf{\hat{e}}^{g}$. At training time, we feed both $T$ and $\hat{T}$ into the text encoder, to obtain their text embeddings $\mathbf{e}^{t}$ and $\mathbf{\hat{e}}^{t}$. In addition to the $\mathcal{L}_{CLIP}$ loss between $\mathbf{e}^{g}$ and $\mathbf{e}^{t}$, we compute ^$\hat{\mathcal{L}}_{CLIP}$ between $\mathbf{\hat{e}}^{g}$ and $\mathbf{\hat{e}}^{t}$. This introduces flexibility in the model to learn an object-centric image embedding in $\mathbf{e}^{g}$, and a more spatially-aware image embedding in $\mathbf{\hat{e}}^{g}$. Both back-propagate into the dense feature maps to learn improved patch embeddings $\{\mathbf{e}_{n}\}_{n=1}^{N}$. At inference time, the model can have access to both types of global image embeddings, and the one to use may be decided based on the downstream task: generally spatially aware tasks will use $\mathbf{\hat{e}}^{g}$ while object-centric ones will employ $\mathbf{e}^{g}$.
@@ 3.2 Integrating Self-Distillation and Masking to Boost Image Features
Discussion. Our method builds on ideas from the weakly and self-supervised literature, and to the best of our knowledge is the first to demonstrate that simultaneously combining contrastive image-text learning with both self-distillation and masked image modeling leads to improvements across many tasks, indicating positive synergies between these objectives. The closest existing techniques are MaskCLIP (Dong et al., 2023) and SILC (Naeem et al., 2024), which combined CLIP with either masked image modeling or self-distillation. As we show in experimental ablations, though, combining the masked image loss with self-distillation substantially improves performance across dense tasks, being critical for downstream applications. We also note some key differences compared to previous methods. Given that we use a CLIP loss, the self-supervised components can be simplified, compared to the original formulations in DINO (Caron et al., 2021) and iBOT (Zhou et al., 2022). A major difference is that we use a single global "crop", instead of two in DINO, iBOT and SILC, substantially increasing throughput by $25\%$. In contrast to many self-supervised methods, we use comparatively simple data augmentations: the local crop is just a random crop of the original image, and the global crop is a larger random crop with horizontal flips. This is similar to Assran et al. (2023); Moutakanni et al. (2024) who argue that complex augmentations may not be necessary for representation learning. Finally, our masking approach is simply random, in contrast to blockwise in iBOT.
@@ 3.3 Scaling TIPS
Data. We leverage the WebLI dataset (Chen et al., 2023), which is a large and noisy web dataset of public images and associated alt-text containing $10$B image-text pairs. We filter the dataset in successive rounds in order to enhance its quality for model training, similar to previous work in language (Gunasekar et al., 2023; Wenzek et al., 2020) and vision (Oquab et al., 2024; Parthasarathy et al., 2023). This is critical for our model, since it is intended for off-the-shelf use in many downstream applications. First, similar to Schuhmann et al. (2022), we filter the image-text pairs based on their contents, by discarding those whose image-text similarities are low, as computed by a pretrained alignment model. Second, we filtered the resulting dataset to only keep pairs with English captions. These two initial steps result in a dataset of $1.7$B images. Finally, we follow a similar curation process as previous work (Oquab et al., 2024; Parthasarathy et al., 2023) and select images that are similar enough to those in curated datasets, leveraging a pretrained model to compute image embeddings; further details on the curated datasets and filtering strategies are reported in Appendix A.3. Note that we also remove near-duplicate images from our dataset if they appeared in any of the evaluation datasets used in this paper. This process generates our main curated pretraining dataset, containing $116$M image-text pairs in total.
@@ 4.1 Experimental Setup
Evaluation datasets and protocols. Our models are evaluated on a suite of $8$ tasks involving $16$ datasets in total, comprising images-only or images-and-text tasks. We assess the quality of the learned representations thoroughly in a wide range of conditions, covering indoor/outdoor scenes and object-centric captures. Note that, in all evaluations, our image-text representations are kept frozen, since our goal is to assess their applicability as off-the-shelf feature extractors. We evaluate $3$ dense prediction tasks, $2$ holistic global image understanding tasks and $3$ multimodal retrieval tasks. We introduce the tasks below, and provide further details about their evaluation protocols in the appendix.
Semantic segmentation is a dense task evaluated on PASCAL VOC (Everingham et al., 2010) and ADE20k (Zhou et al., 2017) datasets, using mean Intersection over Union (mIoU). We use a simple linear probe setup similar to (Oquab et al., 2024), where classes are predicted from the spatial features.
@@
Zero-shot classification is conducted on ImageNet-$1$K by retrieving the class text embedding closest to the each test image’s embedding, following Radford et al. (2021), and using top-1 accuracy.
Additional training dataset details. As discussed in Sec. 3.3, we use images from a set of curated datasets as queries for mining among a large pool of web images. Following DINOv2 (Oquab et al., 2024), we use the training sets of some of our evaluation datasets as the curated queries (details in the appendix). This leads to a web-based training dataset with $116$M image-text pairs, which we use for all released models. Additionally, for the ViT-g models, we experiment with adding the training set of the Mapillary SLS dataset (Warburg et al., 2020) as-is to our training set to compensate for the lack of street-level imagery in web images, and in the absense of any alt-text we use the generated synthetic caption for training both CLS tokens. This increases the number of images in our training set to $117$M in total. A similar procedure is conducted by DINOv2 for their LVD-$142$M dataset.
@@
(Pascal VOC) (NYUv2) (ImageNet) (Flickr) (Flickr)
@@
Implementation details. We use $1$ global crop at resolution $224$ and $M=6$ local crops at resolution $98$. We train the ViT-B models for $70$ epochs at batch size $16$k, which takes $4$ days on $256$ TPUv3 chips. For the ViT-g model we train for $15$ epochs at batch size $16$k, which takes $2$ days on $512$ TPUv5 chips, and results in our low-res model (TIPS-g/14 LR). For our high-res variant (TIPS-g/14 HR), we run an additional finetuning stage with global crops at resolution $448$ and local crops at resolution $140$, for $0.1$ epochs at batch size $4$k. We use only random resize crops and horizontal flips as image augmentations. Models marked with the [rel] prefix correspond to those which are released publicly, whose training set does not include Mapillary SLS. See the appendix for more details.
Compared techniques. We strive to provide a large number of comparisons against recent work. For each existing model family, we compare against the largest instantiation up to ViT sizes “g” or “G”, at about $1.8$B parameters or less in the image encoder. We benchmark TIPS against a wide range of methods, from the self-supervised, weakly-supervised and supervised literature. All methods are used off-the-shelf, with frozen weights, for fair comparisons. As self-supervised methods, we compare against DINO (Caron et al., 2021), MAE (He et al., 2022), iBOT (Zhou et al., 2022) and DINOv2 (Oquab et al., 2024). As weakly-supervised methods, we compare against CLIP (Radford et al., 2021), OpenCLIP (Cherti et al., 2023), SigLIP (Zhai et al., 2023), MaskCLIP (Dong et al., 2023), SILC (Naeem et al., 2024) and EVA-CLIP (Sun et al., 2023). As a supervised method, we benchmark against the ViT-g trained on JFT-$3$B, as per (Zhai et al., 2022).
4.2 Results
@@
[rel] TIPS-g/14 LR (ours) 82.0 47.4 0.390 0.063 23.5 24.7 71.5 83.6 86.3
[rel] TIPS-g/14 HR (ours) 83.1 49.4 0.363 0.059 22.3 24.3 68.4 83.2 86.1
Table 2: Image-only evaluations for dense and global prediction tasks. Experiments using the largest backbone available for each model variant, comparing recent self-supervised and image-text models against TIPS. Rows that are highlighted refer to image-only models that are very good at dense prediction tasks, but are by nature unable to perform text-related tasks. LR (HR) refers to our low-res (high-res) model variant. Models marked with the [rel] prefix refer to those released publicly. We also highlight the best and second-best number of each column. TIPS achieves the best or second-best performance in $7$ out of $9$ evaluations.
@@
Part (B) of the table ablates the contribution of enhanced textual supervision. Simply replacing the web captions by PaliGemma-generated ones improves segmentation by $10.1$ percentage points and reduces depth RMSE by 0.076, which are big positive gains. This shows the potential of synthetic captions for dense understanding with image-text models. However, at the same time global tasks show significant regressions, with KNN classification loss of $6.9$ points. But the CLIP performance can be improved in all tasks by combining the web and synthetic captions: using our dual embedding approach, we achieve large gains across the board. We also compare our dual approach against two other caption combination options, inspired by the ones proposed by Fan et al. (2023): "sampled", where either the web or the synthetic caption is chosen at random; or "multi-text", where both captions are matched against the same image embedding. Our dual approach performs better than other caption combinations in $3$ out $5$ cases and achieves competitive results on the other $2$, which indicates its effectiveness.
@@
[rel] TIPS-g/14 LR (ours) 73.3 93.4 55.6 58.1 82.1 57.0 79.6
[rel] TIPS-g/14 HR (ours) 74.0 93.8 57.0 59.2 83.8 59.4 79.7
Table 3: Image-text evaluations for multimodal retrieval and zero-shot classification, where TIPS outperforms others in $6$ out of $7$ cases. We compare solely against weakly-supervised methods, since self-supervised ones are not naturally aligned with language. We highlight the best and second-best number of each column.
Application: Single-image to 3D. Modern large reconstruction models rely on high-quality pre-trained image encoders to produce image tokens for an encoder/decoder transformer (Hong et al., 2024; Wang et al., 2024). For example, LRM (Hong et al., 2024) predicts parameters of a neural rendering model from the image features of a single input image.
@@
The authors choose the ViT-based DINO encoder over more semantic-aware ones (such as CLIP) due to its knowledge of structural and texture information necessary for 3D tasks. To better understand our model’s capabilities for neural 3D reconstruction, we evaluate TIPS in the LRM framework and compare DINO-B/16 to an equivalently-sized TIPS-B/14. We opt to use DINO-B/16 to follow after the original paper’s implementation. Single-image to $3$D results on the Objaverse (Deitke et al., 2023) dataset are presented in Tab. 4, showing that TIPS outperforms DINO as an image encoder for large reconstruction models, with enhanced novel view synthesis capabilities ($0.62$ increase in PSNR). Qualitative results can be found in the appendix.
@@
Figure 4: Qualitative dense prediction results. For a given image (first column), we illustrate the principal components of the predicted spatial features (column 2) . Depth (column 3) and normals (column 4) are trained on NYUD, and for semantic segmentation (last column) we used the model trained on ADE20k. All dense tasks used the DPT decoder, with a frozen image encoder. More qualitative results can be found in the appendix.